Performance Task #5: Saving Chemistry

Below is the enhanced text of chemistry set problems gathered from the notes of Antoine Laurent Lavoisier during the accident. Please do your best to help by answering the problems and showing your solutions.

I- Part : Concept of Mole

1.      A chemist has a jar containing 388.2 g of iron filings. How many moles of iron does the jar contain?

2.      A student needs 0.366 mol of zinc for a reaction. What mass of zinc in grams should the student obtain?

3.      Calculate the mass of aluminum that would have the same number of atoms as 6.35 g of cadmium.

4.      The weight of a diamond is given in carats. One carat is equivalent to 200. mg. A pure diamond is made up entirely of carbon atoms. How many carbon atoms make up a 1.00 carat diamond?

5.      Precious metals are commonly measured in troy ounces. A troy ounce is equivalent to 31.1 g. How many moles are in a troy ounce of gold? How many moles are in a troy ounce of platinum? of silver?


1
Expert's answer
2022-06-10T08:16:04-0400

1.

M (Fe) = 56 g/mol

m (Fe) = 388,2 g

n = ?


n = m / M = 388,2 g / 56 g/mol = 6,932 mol

Answer: 6,932 mol Fe


2.

M (Zn) = 65 g/mol

n (Zn) = 0,366 mol

m = ?


m = n * M = 0,366 mol * 65 g/mol = 23,79 g

Amswer: 23,79 g Zn


3.

M (Cd) = 112,4 g/mol

M (Al) = 27 g/mol

m (Cd) = 6,35 g

n (Cd) = n (Al)

m (Al) = ?


a).

n (Cd) = m (Cd) / M (Cd)

n(Cd) = 6,35 / 112,4 = 0,0565 mol

b).

n(Cd) = n(Al) = 0,0565 mol

c).

m = n * M

m(Al) = 0,0565 mol * 27 g/mol = 1,5255 g

Answer: 1,5255 g Al


4.

1 carat diamond = 200 mg = 0,2 g

M (C) = 12 g/mol


n = m / M = 0,2 / 12 = 0,01667 mol C


According to Avogadro's law, 1 mole of any substance contains 6.02 * 10 ^ 23 particles (molecules, atoms).

1 mol C — 6,02*1023 atom

0,01667mol C — X

X = 1 * 1022 atom

Answer: 1*1022 atoms


5.

1 troy ounce = 31,1 g

M (Au) = 197 g/mol

M (Pt) = 195 g/mol

M (Ag) = 108 g/mol


n = m / M


n(Au) = 31,1 / 197 = 0,158 mol gold

n(Pt) = 31,1 / 195 = 0,1595 mol platinum

n(Ag) = 31,1 / 108 = 0,288 mol silver